OpenStudy (ttop0816):

please help! x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of _____. -1 1 3

OpenStudy (jojo4eva):

Your answer would be -1

OpenStudy (jojo4eva):

What you do is for the second equation subtract 2x from both sides and u get y=-2x+5 then you substitute that in for y of the first equation and you should have x+5(-2x+5)=-2. Use the distributive property and and u should get x+(-10x)+25=-2. So when you compare like terms you get -9x+25=-2. subtract 25 from both sides and u get -9x=-27 divide both sides by -9 and x=3. then substitute that 3 in for x of the second equation so u can find y. so it would be y=-2(3)+5= -6+5=-1 so y=-1

OpenStudy (jojo4eva):

Any questions? If not then if you need help with anything else let me know

OpenStudy (jojo4eva):

and btw i fanned you

OpenStudy (tester97):

great explanation :) @jojo4eva

OpenStudy (ttop0816):

@jojo4eva thank you so much for your perfect explanation!

OpenStudy (ttop0816):

would you help me on this question too? 7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown? 28x - 12y = 16 and -6x + 12y = -3 14x - 6y = 4 and -14x + 28y = 1 -28x + 12y = -16 and 28x - 56y = 14

OpenStudy (jojo4eva):

Ya

OpenStudy (jojo4eva):

I believe the answer is A please someone correct me if im wrong. THe reason why i say is that if you look at A instead of cancelling out the X's in order to solve for y you are cancelling out the Y's whereas in #s 2 and 3 you are canceling out the X's

OpenStudy (jojo4eva):

@tester97 please correct me if im wrong

OpenStudy (anonymous):

Alternatively, multiply the first equation by -2. Add the two equations together to eliminate the 2x terms. Get -9y = 9, y=-1.

OpenStudy (jojo4eva):

Yes he is correct

OpenStudy (jojo4eva):

@tester97 am i wrong on the second one?

OpenStudy (tester97):

yea dats right cos 7x - 3y = 4 times 3 is 28x - 12y = 16 and 2x - 4y = 1 times 3 is -6x + 12y = -3

OpenStudy (jojo4eva):

Ok so anything else @ttop0816

OpenStudy (ttop0816):

@jojo4eva i promise this is the last one! x + y = k x - y = k The solution to the system shown is _____. (0, k) (2, 2) (k, 0) thank you so much for @jojo4eva & @tester97 )): if it wasnt for you guys i wouldnt have understood these problems!

OpenStudy (jojo4eva):

I believe the answer is B. @tester97 Correct me if im wrong i want to make sure she has the correct answer before she goes.

OpenStudy (tester97):

well the y cancels it self out so that leaves you with x=k both times. so we now know what x is correct?

OpenStudy (jojo4eva):

tru tru

OpenStudy (tester97):

so we can eliminate answer choice A and B right?

OpenStudy (jojo4eva):

got it got it!!! Do u get it @ttop0816